# Adaptive Feature Selection for No-Reference Image Quality Assessment by Mitigating Semantic Noise Sensitivity

### v2

@@ 4.6 Qualitative Analysis
Analyze similar features that have different quality.  In Fig. 4, sections (a) and (b) display the number of image pairs with the most similar features within a specific label distance range in the dataset, as determined by the TReS and QFM-IQM models. The vertical axis in these graphs represents the number, while the horizontal axis shows the range of label differences. Specifically, in Fig. 4(a), it’s noted that out of 1162 distorted images in the LIVEC dataset, 52 images exhibit similar features that have a considerable quality score difference of over 30. This finding highlights a common issue where similar features do not necessarily correspond to similar quality scores. This discrepancy could be due to the presence of only a few quality-relevant features influencing the model’s evaluation. Moreover, Fig. 4(c) conducts a summary statistic of (a) and (b) and demonstrates the effectiveness of our proposed model in addressing this issue. The QFM-IQM model achieves a notable reduction, approximately 56%, in the number of similar features with significant score differences when compared to the TReS model. Additionally, there’s an increase of about 16.3% in the number of similar features having closer scores. This improvement leads to a better positive correlation and, consequently, enhances the overall quality prediction performance of the model. In summary, our model prioritizes quality-related features more than the TReS. This approach enhances the inclusion of these features in the extracted feature set and significantly reduces bias in quality estimation.
Visualization of attention map. We utilize GradCAM (Selvaraju et al., 2017) to visualize the feature attention maps in Fig. 5. The results show that our model accurately and comprehensively focuses on the distorted regions (including both quality-related semantic degradations and low-level artifacts), while DEIQT (Qin et al., 2023) fails to accurately pinpoint the true areas of quality degradation. This indicates that our model excels in learning the subtle differences between semantic structures and quality-aware features. We achieve this by filtering out features irrelevant to quality from the pre-trained model, focusing on learning those quality-aware features. It is worth noting that for generally blurry images, such as those in the third column, our model tends to focus on subject areas that are inherently more prone to becoming blurry, such as floating balloons.
@@ 4.8 Sensitivity study of hyperparameters
In this paper, we use ${\lambda_{1}}$ and ${\lambda_{2}}$ in Eq. 2 to balance the contrastive learning and distillation, respectively. To this end, we conduct a sensitivity study on different feature perturbation weights ${\lambda_{1}}$ and ${\lambda_{2}}$ to explore the effect of QCC, maintaining the search number K at 1. As shown in Table 7 and Table 8, our findings reveal that the QCC module has relatively low sensitivity to the hyperparameters. Specifically, small values of $\lambda$ weaken our QCC’s impact, whereas large values of $\lambda$ cause significant changes in the feature space and result in performance degradation.
@@ Appendix C Training and Evaluation Details
Implement details of training phrase: To train the student network, we adopt a standard approach of randomly cropping input images into ten patches, each with a $224\times 224$ resolution. Subsequently, we reshape these patches into a sequence of smaller patches with a patch size of p = 16 and an input token dimension of D = 384. Furthermore, we present additional training preprocessing details for various datasets, which are not included in the main paper, in Tab. 9. For different benchmarks, we employ different training settings for a fair comparison.
@@ Appendix D Comparison with CONTRIQUE
Our approach differs significantly from CONTRIQUE. In CONTRIQUE, the approach revolves around treating the IQA representation learning problem as a classification task. In this framework, images are categorized into various classes based on distortion types, and contrastive loss is utilized to learn distinctive features that can discriminate between these classes. However, a limitation of this approach is that each real image is treated as a separate class, which doesn’t allow for the effective mining of information from distorted images with similar quality. In contrast, our method does not fixate on specific distortion types. Instead, it takes into account differences in quality scores, avoiding the aforementioned limitation. We achieve this by comparing images and selecting those with similar quality scores but varying semantic features. These selected images are then fused with the original features as a form of noise, to diminish their semantic importance. This approach effectively separates the attributes related to image quality from the semantic features. As shown in Tab.12 and Tab.11), our QFM-IQM shows better performance on performance comparison and cross-dataset validation experiments.

### v3

@@ 4.6 Qualitative Analysis
Analyze similar features that have different quality. In this section, we illustrate our motivation by counting the number of most similar features extracted by the TReS and QFM-IQM models from image pairs with different quality score differences. As shown in Fig. 4, the vertical axis represents quantity, and the horizontal axis represents the range of label differences. According to Fig. 4(a), we observe that SOTA pre-trained models like TReS struggle to distinguish feature distances between image pairs with different quality scores. Specifically, in Fig. 4(a), TReS incorrectly identifies 52 image pairs with a quality score difference greater than 30 as having very similar features. In this paper, We believe this is due to feature confusion, where the model is misled by similar semantics, causing it to perceive similar features in images with different quality scores. This leads to incorrect quality predictions. Our approach mitigates this issue by reducing the model’s sensitivity to semantic noise. Fig. 4(c) summarizes (a) and (b), showing our model’s effectiveness. Compared to TReS, the QFM-IQM model reduces the number of image pairs with similar features but different scores by approximately 56%, while increasing pairs with close scores and similar features by about 16%. This indicates a better positive correlation between features and quality scores, enhancing overall prediction performance. Additionally, QFM-IQM shows fewer high prediction errors (fewer instances in red), further proving its effectiveness. In conclusion, our model emphasizes quality-related features more than TReS, enhancing quality-aware features and significantly reducing quality estimation bias.
Visualization of attention map. We utilize GradCAM (Selvaraju et al., 2017) to visualize the feature attention maps in Fig. 5. The results show that our model accurately and comprehensively focuses on the distorted regions (including both quality-related semantic degradations and low-level artifacts), while DEIQT (Qin et al., 2023) fails to accurately pinpoint the true areas of quality degradation. This indicates that our model excels in learning the subtle differences between semantic structures and quality-aware features. We achieve this by filtering out features irrelevant to quality from the pre-trained model, focusing on learning those quality-aware features. It is worth noting that for overall blurry images, such as those in the third column, our model tends to focus on subject areas that are inherently more prone to becoming blurry, such as floating balloons.
@@ 4.8 Sensitivity study of hyperparameters
In this paper, we use ${\lambda_{1}}$ and ${\lambda_{2}}$ in Eq. 2 to balance the contrastive learning and distillation, respectively. To this end, we conduct a sensitivity study on different feature perturbation weights ${\lambda_{1}}$ and ${\lambda_{2}}$ to explore the effect of QCC, maintaining the search number K at 1. As shown in Table 7 and Table 8, our findings reveal that the QCC module has relative sensitivity to the hyperparameters. Specifically, small values of $\lambda$ weaken our QCC’s impact, whereas large values of $\lambda$ cause significant changes in the feature space and result in performance degradation.
@@ Appendix C Training and Evaluation Details
Implement details of training phrase: To train the student network, we adopt a standard approach of randomly cropping input images into ten patches, each with a $224\times 224$ resolution. Subsequently, we reshape these patches into a sequence of smaller patches with a patch size of p = 16 and an input token dimension of D = 384. Furthermore, we present additional training preprocessing details for various datasets, which are not included in the main paper, in Table 9. For different benchmarks, we employ different training settings for a fair comparison.
@@ Appendix D Comparison with CONTRIQUE
In CONTRIQUE, the approach revolves around treating the IQA representation learning problem as a classification task. In this framework, images are categorized into various classes based on distortion types, and contrastive loss is utilized to learn distinctive features that can discriminate between these classes. However, a limitation of this approach is that each real image is treated as a separate class, which limits the effective mining of information from distorted images with similar quality. In contrast, our QFM-IQM is not confined to a specific type of distortion. Instead, we approach learning quality-aware features from a different perspective, aiming to reduce the sensitivity of the pre-trained network to semantic content through adversarial learning. This technique effectively separates the attributes related to image quality from the semantic features. As shown in Table 12 and Table 11, QFM-IQM exhibits excellent performance in both comparative and cross-dataset validation experiments.